# SandCell: Sandboxing Rust Beyond Unsafe Code

### v1

@@ 9 Related Work
. Several studies have investigated the compartmentalization of applications based on different underlying primitives, including hardware-assisted in-process isolation [54, 41, 36, 60, 52, 58, 43, 57, 63, 61, 40, 38, 22, 35], process-based isolation [42], and software-based isolation [59]. Most of these approaches focus on compartmentalizing applications written in C or C++ [41, 36, 60, 52, 58, 43, 57, 63, 61, 38, 22, 35]. Additionally, several studies specifically address isolating unsafe code from safe code for Rust application [18, 53, 39, 48, 20]. However, those approaches implement conservative compartmentalization policies that label any objects potentially used by untrusted code as unsafe. This can result in edge cases where most allocations are relegated to the unsafe domain, which in turn diminishes the intended benefits of compartmentalization. Also, SDRaD-FFI [34] and SandCrust [42] focus on providing a finer-grained isolation boundary with manual function annotation, but they do not scale effectively for larger isolation boundaries, such as an entire Rust crate. In contrast, SandCell provides a highly automated solution for sandboxing both safe and unsafe code in Rust applications.
@@
. One idea of language-based isolation is to provide language constructs that allow developers to specify isolation boundaries and manage sandboxes. For example, Enclosure [32] allows developers to specify capabilities of libraries. GOTEE [33] puts a Go routine into an Intel SGX enclave and automatically extracts sandboxed code and data from the original code. Both systems allow the sandbox boundary based on one existing language construct (libraries or Go Routines). SandCell further pushes this direction by providing flexible boundaries such as crates, functions and types. Also, prior work target Go applications, while SandCell targets Rust applications. Another idea is to provide a sandbox interface in the language or through library, to support sandboxed execution of external code. JavaScript engine V8 [15] can be used as a virtual machine that runs user code, whose interaction with the outside world is managed by V8. WebAssembly [16] provides a sandboxed execution environment for web applications by including bounds checking in the semantics of the instruction set. Libsandbox [3] provides APIs in C/C++ and Python for executing and profiling single process programs in a sandbox. However, these tools are designed to execute foreign code in a one-off manner with a high cost of context switching. Furthermore, retrofitting legacy code into these APIs is troublesome. In contrast, SandCell leverages existing syntax boundaries to alleviate the effort needed for compartmentalization.
@@
10 Discussion & Future Work
The current prototype of SandCell has the following limitations. First, in MIR, the functions are not monomorphized, so if at a call site, the callee function’s concrete type cannot be determined by the call graph construction, the instrumentor will not be able to generate the copying code for the argument with a generic type that might have different sizes at runtime. Second, in Rust, a dynamic trait object is implemented as a fat pointer that contains a pointer to the object and a pointer to the vtable. Without knowing the concrete type of the object, the analysis cannot trace down the references contained in the object. Both issues can be resolved by doing instrumentation at the LLVM IR level and we leave them as future work. In our evaluation, we encountered Box<dyn Trait> types crossing the boundary, which is a heap pointer pointing to a virtual table and an opague object. The type of the object is not known at compile time, which would require reflection at runtime to determine how to copy the object across the boundary. However, Box itself is a heap object, so allocating it in the shared domain can solve the problem if the object does not contain references to other object.
@@ 10 Discussion & Future Work
Open Science
Appendix
Task SandCrust SDRaD-FFI SandCell
arithmetic 100.89% 8.31% -0.06%
my_blood 106.10% 7.66% 1.38%
objects 66.75% 5.53% -0.50%
rust_logo 81.75% 8.95% 1.27%
Table 3: The Percentage of Overhead for Sandboxing libpng Compared to Baseline (No Sandboxing).
Matrix Size Baseline SandCell-Copy Overhead SandCell-Share Overhead
@@
Table 4: Running Time (milliseconds) of RustFFT with transpose Sandboxed on Square Matrices of Different Lengths.
File Numbers Baseline SandCell-Share Overhead
@@
[…]
Table 5: ripgrep’s Average Running Time and Standard Deviation on Different Number of Files.
File Size Baseline SandCell-Copy Overhead of SandCell-Copy SandCell-Share Overhead of SandCell-Share
47 8.193 $\pm$ 2.857 ms 9.234 $\pm$ 3.269 ms 12.7% 9.817 $\pm$ 3.321 ms 19.8%
195 8.393 $\pm$ 3.199 ms 7.765 $\pm$ 2.792 ms -7.5% 10.412 $\pm$ 3.539 ms 24.1%
1192 7.850 $\pm$ 3.109 ms 8.536 $\pm$ 2.972 ms 8.7% 8.925 $\pm$ 3.564 ms 13.7%
Table 6: The Running Time (seconds) of Parsing ELF Files with elf-rs, shown as Mean and Standard Deviation ($\sigma$).
Connection Baseline SandCell-Copy SandCell-Share
2 92.7 21742.86 109.59 19219.25 102.32 19490.26
4 109.69 37639.96 137.27 32279.09 153.62 31055.12
6 135.41 48953.95 146.62 36015.92 166.39 41210.53
8 148.75 47132.76 176.62 47790.42 178.13 48427.64
10 191.83 61150.93 190.13 51648.45 209.21 52390.42
Table 7: rouille Performance with Different Numbers of Concurrent Connections.
Task Baseline SandCell-Copy Overhead (%) SandCell-Share Overhead (%)
[…]
Table 8: The Running Time (milliseconds) of an HTTP Client backed by openssl on Different Tasks.

### v2

@@ 9. Related Work
. Several studies have investigated the compartmentalization of applications based on different underlying primitives, including hardware-assisted in-process isolation (Rivera16, ; Koning17, ; Hedayati19, ; Vahldiek-Oberwagner19, ; Melara19, ; Sung20, ; Lefeuvre21, ; Schrammel20, ; Wang20, ; Voulimeneas22, ; Kirth22, ; Jin22, ; Chen22, ; gulmez2023dsn, ), process-based isolation (lamowskiSandcrustAutomaticSandboxing2017, ), and software-based isolation (Tan17, ). Most of these approaches focus on compartmentalizing applications written in C or C++ (Koning17, ; Hedayati19, ; Vahldiek-Oberwagner19, ; Melara19, ; Sung20, ; Lefeuvre21, ; Schrammel20, ; Wang20, ; Voulimeneas22, ; Jin22, ; Chen22, ; gulmez2023dsn, ). Additionally, several studies specifically address isolating unsafe code from safe code for Rust application (almohriFideliusCharmIsolating2018, ; riveraKeepingSafeRust2021, ; kirthPKRUsafeAutomaticallyLocking2022, ; liuSecuringUnsafeRust2020, ; bangTRustCompilationFramework2023, ). However, those approaches implement conservative compartmentalization policies that label any objects potentially used by untrusted code as unsafe. This can result in edge cases where most allocations are relegated to the unsafe domain, which in turn diminishes the intended benefits of compartmentalization. Also, SDRaD-FFI (gulmezFriendFoeExploring2023, ) and Sandcrust (lamowskiSandcrustAutomaticSandboxing2017, ) focus on providing a finer-grained isolation boundary with manual function annotation, but they do not scale effectively for larger isolation boundaries, such as an entire Rust crate. In contrast, SandCell provides a highly automated solution for sandboxing both safe and unsafe code in Rust applications.
@@
. One idea of language-based isolation is to provide language constructs that allow developers to specify isolation boundaries and manage sandboxes. For example, Enclosure (ghosnEnclosureLanguagebasedRestriction2021, ) allows developers to specify capabilities of libraries. GOTEE (ghosn2019secured, ) puts a Go routine into an Intel SGX enclave and automatically extracts sandboxed code and data from the original code. Both systems allow the sandbox boundary based on one existing language construct (libraries or Go Routines). SandCell further pushes this direction by providing flexible boundaries such as crates, functions and types. Also, prior work target Go applications, while SandCell targets Rust applications. Another idea is to provide a sandbox interface in the language or through library, to support sandboxed execution of external code. JavaScript engine V8 (GoogleV8, ) can be used as a virtual machine that runs user code, whose interaction with the outside world is managed by V8. WebAssembly (WebAssembly, ) provides a sandboxed execution environment for web applications by including bounds checking in the semantics of the instruction set. Libsandbox (githubGitHubOpenJudgeSandbox, ) provides APIs in C/C++ and Python for executing and profiling single process programs in a sandbox. However, these tools are designed to execute foreign code in a one-off manner with a high cost of context switching. Furthermore, retrofitting legacy code into these APIs is not desirable for large-code bases. In contrast, SandCell leverages existing syntax boundaries to alleviate the effort needed for compartmentalization.
@@
10. Discussion & Future Work
In-process isolation.
Like other in-process isolation relying on Intel MPK, SandCell is constrained by the hardware limit of 16 distinct protection domains. As SandCell’s specification and analysis can work independently of the underlying enforcement mechanism, future work could explore replacing MPK with alternative in-process isolation schemes, such as CHERI-based hardware capabilities (cheri, ) or SFI (Tan17, ).
System Call Filtering Alternatives
. As we discuss in Section 7.4, SandCell integrates the zpoline system call interpositioning as part of SDRaD-SandCell to harden PKU-based isolation and to address Challenge IV. While SDRaD-SandCell is fully compatible with conventional system call filtering techniques used in prior PKU-based isolation approaches (Vahldiek-Oberwagner19, ; Voulimeneas22, ), these techniques typically require kernel modifications to improve their efficiency. Without kernel support, they tend to incur significant performance overhead.
System Call Filtering Enhancements
. Currently, SandCell’s isolation only focuses on memory isolation. However, its system call interpositioning mechanism could also be used to enforce more system call policies. For example, we can extend the specification to allow developers to deny certain system call for a particular sandbox. We leave this as future work.
11. Conclusion
We presented SandCell, a novel compiler-based tool that analyzes and instruments Rust code to provide component isolation; we believe that our open-source release of SandCell paves the way for future advancements in securing the Rust ecosystem.
@@ 11. Conclusion
Generative AI Usage
We used OpenAI Codex to assist in the debugging process of our Rust codebase and to generate our evaluation scripts/workloads. The debugging output can be straightforwardly verified. The evaluation scripts/workloads were reviewed and modified by the authors to ensure correctness. We also use Google’s Gemini to help searching for related work and benchmarks. The search results were further examined by the authors to ensure relevance. For grammar and spelling checking, we used GitHub Copilot.
Appendix A Open Science
The artifacts are available at https://anonymous.4open.science/r/SandCell-Artifact-2D4E. The artifacts include the implementation of the tool, SandCell, and all the benchmarks used in the evaluation, along with scripts to reproduce the results. The instructions for reproducing the results are included in the README file in the artifact repository.
Appendix B Supplementary Measurement
We list all the data from the evaluation below. For readability, we include both the plots and tables for these experiments.
((a)) Running Time of elf-rs
((b)) Running Time of transpose
((c)) Running Time of ripgrep
((d)) Running Time of async-graphql
((e)) Running Time of oq
((f)) Running Time of attohttpc
Figure 6. Performance benchmarks.
Image Baseline SandCrust SandCell Overhead of Overhead of Overhead
Size SandCell SDRaD-FFI SDRaD-FFI
5.5K $160.17$ +183.06% $164.30$ +2.58% $176.28$ +10.06%
63K $2391.45$ +225.99% $2389.73$ -0.07% $2506.35$ +4.80%
378K $5797.16$ +64.21% $5844.68$ +0.82% $5904.48$ +1.85%
874K $8961.68$ +129.44% $8943.76$ -0.20% $9306.31$ +3.85%
Table 3. The Percentage of Overhead for Sandboxing libpng Compared to Baseline.
Matrix Size Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
64 $512.370(22.420\sigma)$ $537.210(21.470\sigma)$ $4.8\%$ $535.010(22.300\sigma)$ $4.4\%$
@@ Appendix B Supplementary Measurement
96 $1476.160(38.400\sigma)$ $1520.240(36.120\sigma)$ $3.0\%$ $1525.440(54.950\sigma)$ $3.3\%$
112 $1892.640(65.620\sigma)$ $1961.020(45.780\sigma)$ $3.6\%$ $1954.820(64.460\sigma)$ $3.3\%$
128 $2046.130(44.300\sigma)$ $2125.820(57.530\sigma)$ $3.9\%$ $2130.330(45.710\sigma)$ $4.1\%$
Table 4. RustFFT with transpose: Running Time (milliseconds) on Square Matrices of Different Lengths.
Task Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
@@
post $51.061(5.741\sigma)$ $62.377(21.081\sigma)$ $22.2\%$ $60.089(13.794\sigma)$ $17.7\%$
session $182.633(21.552\sigma)$ $195.319(56.078\sigma)$ $6.9\%$ $200.248(33.824\sigma)$ $9.6\%$
multipart $61.411(4.607\sigma)$ $74.763(12.469\sigma)$ $21.7\%$ $72.036(17.354\sigma)$ $17.3\%$
@@
stream $54.910(8.434\sigma)$ $61.294(14.557\sigma)$ $11.6\%$ $62.426(17.467\sigma)$ $13.7\%$
Table 5. HTTP Client backed by openssl: Running Time on Different Tasks.
File Count (8MB each) Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
64 $119.476(11.456\sigma)$ $131.738(13.340\sigma)$ $10.3\%$ $135.476(9.656\sigma)$ $13.4\%$
128 $23.320(12.575\sigma) 22$$760(3.558\sigma) 2.92.9\%$ $233.460(1.739\sigma)$ $4.5\%$
@@
512 $833.026(35.350\sigma)$ $857.149(40.615\sigma)$ $2.9\%$ $874.843(42.189\sigma)$ $5.0\%$
Table 6. ripgrep’s Running Time on Different Number of Files.
File Size Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
47 $9.532(1.680\sigma)$ $0.601(1.442\sigma) 11$$1.2\%$ $10.928(1.347\sigma)$ $14.6\%$
@@
1192 $22.910(4.491\sigma)$ $24.441(3.216\sigma)$ $6.7\%$ $33.654(2.131\sigma)$ $46.9\%$
5805 $61.925(2.556\sigma)$ $71.721(2.248\sigma)$ $15.8\%$ $121.957(2.723\sigma)$ $96.9\%$
Table 7. Parsing ELF Files with elf-rs Running Time
Pool Size Baseline SandCell-Share SandCell-Copy
1 8820.0 1584.41 10130.0 1577.41 13830.0 1155.43
2 5030.0 3180.91 4380.0 3191.21 7360.0 2171.07
4 2540.0 6302.12 2610.0 6115.68 4500.0 3335.04
Table 8. rouille Performance across Different Thread Pool Sizes.
Query Baseline SandCell-Copy Overhead of SandCell-Copy SandCell-Share Overhead of SandCell-Share
static.graphql $34.173(3.280\sigma)$ $39.568(5.655\sigma)$ $15.8\%$ $38.572(2.971\sigma)$ $12.9\%$
dashboard.graphql $29.245(0.882\sigma)$ $33.661(3.156\sigma)$ $15.1\%$ $32(0.878\sigma) 13.313.3\%$
batch.graphql $30.781(2.883\sigma)$ $34.585(1.793\sigma)$ $12.4\%$ $34.121(0.750\sigma)$ $10.8\%$
@@
Table 9. async-graphql Performance Benchmark across Different Query Workloads.
Input Baseline SandCell-Share Overhead of SandCell-Share SandCell-Copy Overhead of SandCell-Copy
100k $13.143(2.519\sigma)$ $17.511(2.717\sigma)$ $33.2\%$ $17.871(3.466\sigma)$ $36.0\%$
200k $22.894(3.975\sigma)$ $24.826(2.458\sigma)$ $8.4\%$ $24.206(1.069\sigma)$ $5.7\%$
336k $33.901(5.077\sigma)$ $37.675(3.065\sigma)$ $11.1\%$ $37.424(3.467\sigma)$ $10.4\%$
400k $37.334(2.568\sigma)$ $41.679(2.327\sigma)$ $11.6\%$ $40.845(3.113\sigma)$ $9.4\%$
500k $45.952(7.905\sigma)$ $50.595(3.499\sigma)$ $10.1\%$ $50.986(3.096\sigma)$ $11.0\%$
Table 10. oq Performance Benchmark for Different Input Size.
Benchmark Overhead of SandCell-Share
Sunspider -0.007%
v8 1.650%
jzlib -1.432%
dom -0.131%
dromaeo 9.360%
Table 11. Servo Average Overhead on each Testsuite of the Dromaeo Benchmark.